OpenStudy (freemap):

The function s(m) = 5 + 4(m – 1) represents the number of stamps in Rudy’s stamp collection in month m. What does the value 4 represent in this situation? Rudy’s collection is worth $4. Rudy collected stamps for 4 months. Rudy began his collection with 4 stamps. Rudy adds 4 more stamps to his collection every month.

OpenStudy (freemap):

I gave you medal for last question. I forgot about it. sorry

Directrix (directrix):

No worries. We're on to the stamp problem now.

Directrix (directrix):

If m = 1, Rudy has 5 + 4*(1-1) = 5 + 0 = 5 stamps.

OpenStudy (freemap):

D?

Directrix (directrix):

I think so. In month 2, he has 5 + 4*(2-1) = 5 + 4 = 9 stamps.

OpenStudy (freemap):

Thank you

Directrix (directrix):

The equation s(m) = 5 + 4(m – 1) simplifies to s(m) = 4m + 1. That slope of 4 means the rate of change is 4/1 or 4 stamps more per 1 month more. I vote for this: Rudy adds 4 more stamps to his collection every month.

Directrix (directrix):

You are welcome. (Is it me or is every answer D)

OpenStudy (freemap):

lol